OpenStudy (anonymous):

what is an explicit rule for the nth term of the sequence for 9, 18, 36, 72

OpenStudy (anonymous):

Let's say the sequence starts with \(n=1\). The the first term (\(n=1\)) is \(9\times2^{1-1}=9\times1=9\), the second (\(n=2\)) is \(9\times2^{2-1}=9\times2=18\), the third (\(n=3\)) is \(9\times2^{3-1}=9\times4=36\), and so on. See a pattern?